# Light of Normals: Unified Feature Representation for Universal Photometric Stereo

### v2

@@ 3.2 PS-Verse Dataset
[27] pioneers the design of a large-scale synthetic dataset, PS-Wild, for Universal PS tasks. It utilizes single-material 3D models from the commercial AdobeStock collection [1], rendering over 10,000 scenes augmented through rotation and color variations. However, the lighting setup is limited to HDRI environment maps, lacking high-frequency point light sources that provide richer illumination effects. Building on this dataset, [31] extends the lighting configuration by combining area lights, environment lights, and point lights to better simulate natural illumination. Additionally, it replaces single objects with multi-object compositions to disrupt semantic cues and enhance model generalization. Nonetheless, these objects lack fine-grained surface normals and complex geometric structures, resulting in relatively limited lighting variations and self-shadowing effects, which restrict further improvements in model performance.
Using the availability of publicly available large-scale 3D asset datasets [12, 11, 51], we propose the PS-Verse dataset, which naturally expands the training data required by Universal PS. To increase the lighting complexity in rendered scenes, we carefully select 17,805 textured 3D models with UV coordinates and PBR materials from the Objaverse dataset [12], which contains nearly 800,000 models. Following Dora [7], which quantifies geometric complexity based on the density of sharp edges, we categorize the objects into four complexity levels. We then construct scenes by recursively selecting 4 to 6 objects from each level, and classify scenes into four corresponding complexity tiers.
Regarding materials, textures are randomly sampled from the MatSynth large-scale PBR material database [51], consisting of 806 metallic, 1,226 specular, and 3,321 diffuse material groups. To balance the distribution of specular and diffuse lighting effects, we assign object materials according to the ratio of 1:4:2.5:2.5 across original object textures, diffuse, specular, and metallic materials. To more realistically simulate fine surface detail interactions with light, we introduce normal maps for objects for the first time in such data, defining this as a fifth complexity level that injects rich high-frequency lighting details.
In terms of lighting setup, we incorporate a uniform environment light to simulate non-dark ambient conditions. Detailed lighting configurations are provided in the appendix. Overall, the PS-Verse dataset consists of 100,000 scenes generated using the Cycles renderer in Blender [15]. Each scene includes two rendering outputs: with and without normal maps. We render 20 16-bit images at a resolution of 512x512 per scene. Specifically, for quantitative metric assessment, we chose 10 scenes from each level (Level 1 to Level 5), collectively constituting what we term our PS-Verse Testdata.
@@ 4 Experiments
Table 2: Ablation results on architecture trained on PS-Mix [31].
@@
SDM-UniPS 0.84 0.72 5.80 13.50 14.96
[…]
Implementation Details: To maintain a data volume comparable to that of PS-Mix [31], we limited our selection from the PS-Verse dataset to 40K scenes for training (to align with PS-Mix [31], which is about 35K). We start training on PS-Verse Level 1 data, adding higher levels every 10 epochs up to Level 4, for about 150 epochs without ground truth normals. Then, we fine-tune on Level 5 data with ground truth normals until 200 epochs to improve fine surface detail reconstruction. We use AdamW Optimizer with $1e^{-4}$ initial learning rate, $0.05$ weight decay and a step decay of 0.8 every ten epochs. Input images per batch vary randomly from 3 to 6. The total loss $\mathcal{L}$ combines multiple components:
$\mathcal{L}=\mathcal{L}_{\text{light}}+\mathcal{L}_{\text{N}}$
Training runs on eight NVIDIA H100 GPUs for roughly seven days. Inference takes around 2 seconds for 16 input images at 512×512 on H100.
Evaluation Metric: We evaluates the results with three metrics. Normal accuracy is measured by Mean Angular Error (MAE), with lower values indicating better performance. Feature similarity is assessed by Cosine Similarity (CSIM) and Structural Similarity Index (SSIM) [54]: encoder-extracted features are normalized, CSIM is computed by cosine similarity, and SSIM is calculated after PCA reduces features to three dimensions. Higher CSIM and SSIM indicate better similarity.
[…]
Table 3: Main ablation with 20 multi-lights input images. The evaluation metrics were measured on our PS-Verse Testdata 3.2.
[…]
Level 1 Level 2 Level 3 Level 4 Level 5 Avg.
SDM-UniPS [31] PS-Mix [31] 0.61 0.63 9.27 10.34 12.64 13.89 14.87 12.02
SDM-UniPS [31] PS-Verse 0.68 0.65 7.35 8.15 8.12 9.47 11.69 8.82
Baseline PS-Verse 0.71 0.69 7.52 8.03 8.41 9.12 10.55 8.73
w/ tokens PS-Verse 0.74 0.73 7.01 8.15 8.34 8.43 8.65 8.13(0.60 $\downarrow$)
w/ global attention PS-Verse 0.80 0.78 5.64 6.41 7.52 7.12 6.51 6.44 (2.29 $\downarrow$)
w/ align PS-Verse 0.86 0.82 4.35 5.60 5.47 6.10 5.97 5.58 (3.15 $\downarrow$)
w/ wavelet transform PS-Verse 0.85 0.82 4.32 5.10 4.96 5.71 5.64 5.15 (3.58 $\downarrow$)
w/ grad perception loss PS-Verse 0.86 0.83 4.18 4.90 4.77 5.54 5.21 4.84 (3.89 $\downarrow$)
[…]
Table 4: Evaluation on LUCES[39]. Uses all 52 images unless otherwise noted (K).
UniPS [28] 11.01 24.12 23.84 27.90 23.51 28.64 16.24 21.41 35.93 14.53 32.87 28.36 25.36 19.03 23.77
SDM-UniPS [31] 13.30 12.76 8.44 18.58 8.53 19.67 7.25 8.86 26.07 8.30 12.67 15.97 16.01 12.54 13.50
Uni MS-PS [20] 10.29 10.51 6.79 12.57 9.60 13.35 6.27 8.44 25.46 6.10 11.38 15.97 11.37 12.22 11.10
@@
Ours (K=15) 10.23 7.35 8.32 14.73 6.26 7.76 7.01 5.91 22.78 7.02 9.93 9.78 9.70 7.53 9.59
Ours (K=6) 10.84 7.85 8.43 16.41 6.43 8.16 9.24 6.39 24.34 6.38 9.96 11.02 10.49 8.22 10.30
Ours (K=3) 11.55 9.53 10.43 19.01 11.11 8.63 8.83 7.01 28.79 5.44 11.72 13.80 13.95 9.35 12.08
4.1 Ablation Study
To illustrate the effectiveness of our proposed method, LINO-UniPS was trained to convergence (approx. 140 epochs) on the PS-Mix dataset [31].
As presented in Tab. 2, our LINO-UniPS surpasses SDM-UniPS, also trained to convergence on PS-Mix, across both evaluated public datasets. Notably, LINO-UniPS yields higher CSIM and SSIM scores, quantitatively demonstrating that features extracted by our encoder exhibit greater similarity compared to those from SDM-UniPS [31]. It is noteworthy that our LINO-UniPS utilizes a decoder architecture identical to that of SDM-UniPS [31]. This architectural commonality strongly suggests that the observed performance improvements are primarily attributable to our encoder’s enhanced capability to more effectively decouple illumination from geometry. Such effective decoupling fosters stronger normal consistency within the learned features, consequently boosting the final normal reconstruction accuracy and capability.
Figure 7: Our method outperforms other universal PS methods in challenging synthetic scenes with no masks, detailed textures, complex geometry, and high reflectivity.
The main ablation results are presented in Tab. 3. Firstly, by comparing the performance of SDM-UniPS when trained to convergence on the PS-Mix dataset versus our PS-Verse dataset, we observe that training on PS-Verse yields superior feature similarity scores (CSIM and SSIM) as well as better MAE for normal reconstruction. This indicates that training on our PS-Verse dataset, compared to PS-Mix, more effectively enhances the capability to reconstruct normals for fine-grained details. Secondly, we conducted ablations of the various modules within our LINO-UniPS method, starting from a baseline model where all our proposed enhancements are removed. Our initial finding is that by merely incorporating unsupervised light tokens, both the feature similarity metrics and the normal reconstruction results show an improvement. This suggests that these additional tokens can often capture global lighting information even without direct supervision, thereby aiding the disentanglement of normals from illumination. Progressively incorporating our global attention mechanism and lighting alignment strategy leads to more significant improvements in both feature similarity and normal reconstruction performance. This further validates our central conclusion: alleviating the coupling between lighting and normal features effectively enhances the quality of normal reconstruction. Finally, the last two rows in Tab. 3 pertain to the wavelet transform and the gradient perception loss, components we introduced for extracting fine-grained contexts. While these two modules exhibit a comparatively modest contribution to the overall feature similarity metrics, their substantial performance improvements on data with complex geometries, particularly for Level 4 and Level 5 of the PS-Verse dataset, underscore their justification and effectiveness in handling intricate surface details.
Table 5: Evaluation on DiLiGenT[48]. Uses all 96 images unless otherwise noted (K).
Method Ball Bear Buddha Cat Cow Goblet Harvest Pot1 Pot2 Reading Avg.
UniPS [28] 4.90 9.10 19.40 13.00 11.60 24.20 25.20 10.80 9.90 18.80 14.70
SDM-UniPS [31] 1.50 3.60 7.50 5.40 4.50 8.50 10.20 4.70 4.10 8.20 5.80
Uni MS-PS [20] 1.84 3.14 6.04 3.45 3.99 6.49 8.90 4.12 4.70 7.00 4.97
@@ 4.2 Quantitative Results
On the DiLiGenT’s [48] 10 objects, our LiNO-UniPS obtains the largest number of SOTA results on individual objects, while its mean performance metrics also ranks SOTA (results can be seen in Tab. 5). On LUCES [37], a dataset predominantly featuring high-frequency information, our quantitative results are presented in Tab. 4. Compared to all evaluated Universal Photometric Stereo (PS) methods, our LINO-UniPS consistently attains SOTA performance.Remarkably, for the House object, which features particularly intricate geometry, our LINO-UniPS outperforms all other compared methods. Furthermore, our approach demonstrates excellent results even with a very limited number of input images (e.g., $K=3$). We attribute this robustness to our strategy of strengthening Lighting-Normal decoupling; this enables our method to extract features with high normal consistency even from sparse inputs, consequently leading to superior reconstruction quality.
Our LINO-UniPS not only achieves the best performance among all Universal PS methods evaluated but also significantly surpasses representative Calibrated PS [44] and Uncalibrated PS approaches [5]. Specifically, for objects exhibiting difficult material properties (e.g., ACRYLIC) or complex shapes (e.g., PENTAGON), our LINO-UniPS demonstrates markedly superior performance. We also conduct a comparative evaluation of existing Universal Photometric Stereo (PS) methods on PS-Verse Testdata 3.2, with the results (detailed in Fig. 7) clearly demonstrating that our approach significantly surpasses other contemporary Universal PS techniques, especially in these more demanding cases.
Figure 8: High-resolution (>2K) surface normal recovery from real captured images under varying lighting conditions. Top: Canandwood from [28], masked object; bottom: Coins and keyboard from [31], unmasked scene.
Finally, to further underscore the efficacy of our encoder, we conducted an additional experiment. In this setup, we replaced the standard decoder in our LINO-UniPS framework with a significantly simpler Multi-Layer Perceptron (MLP) and then fine-tuned this variant until convergence. Our findings were twofold: First, the consistency of the extracted features was further enhanced, as detailed in Fig. 5. Second, although the reconstruction performance saw a slight degradation compared to LINO-UniPS with its original, more sophisticated decoder, this MLP-decoder variant still markedly outperformed SDM-UniPS (which utilizes a complex encoder architecture), with comparative results presented in Tab. 4. These results align with our central hypothesis that the superior performance of LINO-UniPS is primarily driven by its encoder’s advanced capability to generate highly consistent and well-disentangled features.
@@ 4.3 Qualitative Results
Fig. 1 and Fig. 8 showcase our method’s ability to reconstruct highly detailed and accurate surface normals for diverse real-world objects and scenes, highlighting its strong generalization capabilities. Additional reconstruction results on real-world scenarios can be found in appendix.
@@ 5 Conclusion
In this study, we propose LINO-UniPS, a universal photometric stereo solution. We identify the core challenge of this task as decoupling consistent normal information from spatially varying lighting features, achieving invariance amid variation. Specifically, we utilize a lighting register and a global attention mechanism to capture global lighting and enhance depth-context awareness across multiple images, enabling effective separation of lighting and normal features. To improve the recovery of high-frequency normal details, we replace traditional up/downsampling with wavelet transforms to preserve high-frequency information and introduce a normal gradient-aware loss to boost the model’s sensitivity to details. Finally, we contribute a complex and large-scale photometric stereo dataset, hoping to provide valuable reference for future research.
@@
This supplementary document offers further technical details, demonstrations of the datasets employed, additional insights, and comprehensive results pertaining to our LINO-UniPS.
@@
To provide readers with a more in-depth understanding of our LINO-UniPS network architecture, the method primarily comprises seven key modules:(a) WaveDownSample module, (b) DINOv2-based feature extraction backbone, (c) an Enhanced Light-Normal Contextual Attention Module, (d) a DPT-Based Fusion Module, (e) a WaveUpSample module, (f) a final Decoder and (g) the training loss. An overview of our network architecture is presented in Figure 2. In the subsequent sections, we will provide a detailed account of the network’s structural components and the specific transformations of tensor shapes as data progresses through the model.
A1.1.1 WaveDownSample:
To enable our encoder to extract more fine-grained contexts, we first incorporate the wavelet transform [10], chosen for its ability to separate an image’s high- and low-frequency components [14, 23] while concurrently mitigating losses typically incurred during downsampling [43].
@@
Following SDM-UniPS [31], for each pre-processed input image $I_{f}\in\mathbb{R}^{H\times W\times 3}$, we perform two separate transformations: naive downsampling to obtain $I_{f}^{d}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$ in the image domain, and a wavelet transform to yield its corresponding wavelet domain components $I_{f}^{w}\in\mathbb{R}^{4\times\frac{H}{2}\times\frac{W}{2}\times 3}$, namely $I_{f}^{ll}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$, $I_{f}^{lh}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$, $I_{f}^{hl}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$, and $I_{f}^{hh}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$.
A1.1.2 DINOv2 backbone:
@@
Subsequently, both the downsampled image representation $I_{f}^{d}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$ and the set of wavelet components $I_{f}^{w}$ (comprising $I_{f}^{ll},I_{f}^{lh},I_{f}^{hl},I_{f}^{hh}$, each in $\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$) are individually processed. First, each of these input components is partitioned into a sequence of patch-based tokens. These token sequences are then fed into our feature extraction backbone. To maintain a lean model architecture and leverage strong prior knowledge while minimizing the introduction of excessive additional parameters, we employ the DINOv2-small variant [42] as this backbone. This DINOv2-small backbone is initialized with its publicly available pre-trained weights and is subsequently fine-tuned during our training procedure to extract rich visual representations from these diverse inputs. In our specific implementation, we set the patch size to $P=8$. Consequently, for each input component with spatial dimensions $H/2\times W/2$, the resulting sequence length is $L=\frac{(H/2)\times(W/2)}{P^{2}}$ tokens. Given our use of the DINOv2-small variant, its feature embedding dimension is $D=384$. Following processing by this DINOv2 backbone, we respectively obtain shallow visual feature representations $F_{s,f}^{d}\in\mathbb{R}^{L\times D}\in$ from the downsampled image stream (derived from $I_{f}^{d}$) and $F_{s,f}^{w}\in\mathbb{R}^{4\times L\times D}$ from the wavelet components stream (derived from $I_{f}^{w}$).
A1.1.3 Enhanced Light-Normal Contextual Attention Module:
@@
To achieve a more effective decoupling of lighting and normal features that are subsequently processed by the decoder, we introduce our Enhanced Light-Normal Contextual Attention Module.

### v4

@@ 3.2 PS-Verse Dataset
Prior large-scale synthetic datasets, such as PS-Wild Ikehata (2022) and PS-Mix Ikehata (2023), have advanced the field by enabling data-driven Universal PS. However, they remain limited by either overly simplistic lighting (e.g., lacking high-frequency point sources) or geometrically simple objects that fail to produce complex lighting variations and self-shadowing effects.
Leveraging publicly available large-scale 3D asset datasets Deitke et al. (2022; 2023); Vecchio & Deschaintre (2024), we propose the PS-Verse dataset (see Tab. 1 for a detailed comparison with other Photometric Stereo datasets). To increase the lighting complexity in rendered scenes, we carefully select 17,805 textured 3D models with UV coordinates and PBR materials from the Objaverse dataset Deitke et al. (2022), which contains nearly 800,000 models. Following Dora Chen et al. (2024), which quantifies geometric complexity based on the density of sharp edges, we categorize the objects into four complexity levels. We then construct scenes by recursively selecting 4 to 6 objects from each level, and classify scenes into four corresponding complexity tiers.
Regarding materials, textures are randomly sampled from the MatSynth large-scale PBR material database Vecchio & Deschaintre (2024), consisting of 806 metallic, 1,226 specular, and 3,321 diffuse material groups. To balance the distribution of specular and diffuse lighting effects, we assign object materials according to the ratio of 1:4:2.5:2.5 across original object textures, diffuse, specular, and metallic materials. To more realistically simulate fine surface detail interactions with light, we introduce normal maps for objects for the first time in such data, defining this as a fifth complexity level that injects rich high-frequency lighting details.
In terms of lighting setup, we use uniform environment light to simulate non-dark ambient conditions. Detailed lighting configurations are provided in the appendix A2. Overall, the PS-Verse dataset consists of 100,000 scenes generated with Blender Foundation . Each scene includes two rendering outputs: with and without normal maps. We render 20 images at a resolution of 512 per scene.
In addition to providing ground truth for surface normals, the PS-Verse dataset also offers ground truth for albedo, roughness, and metallic, to support PBR prediction tasks. The PBR prediction results and visual showcase of the PS-Verse dataset can be found in the appendix A1.2 and A1.3.
@@ 4 Experiments
Implementation Details: To effectively enhance our method’s capability for reconstructing fine-grained surface normals, we employ a curriculum learning strategy that progresses from low to high geometric complexity. We start training on PS-Verse Level 1 data, adding higher levels every 10 epochs up to Level 4, for about 150 epochs. Then, we finetune on Level 5 data with ground truth normals until 200 epochs to improve surface detail reconstruction. We use AdamW Optimizer with $1e^{-4}$ initial learning rate, $0.05$ weight decay and a step decay of 0.8 every ten epochs. Input images per batch vary randomly from 3 to 6. The total loss $\mathcal{L}$ combines multiple components:
$\mathcal{L}=\mathcal{L}_{\text{light}}+\mathcal{L}_{\text{n}}$
Training runs on 2 NVIDIA H100 GPUs for roughly 3 days. Inference takes around 1.5 seconds for 16 input images at 512×512 on H100.
Evaluation Metric: We evaluate our results using three metrics. For surface normal accuracy, we measure the Mean Angular Error (MAE, in degrees, ↓). Then, we assess feature similarity via Cosine Similarity (CSIM ↑) on normalized features and Structural Similarity Index Wang et al. (2004) (SSIM ↑) on their 3D PCA projections.
Evaluation Dataset: For evaluation, Comparison experiments and ablation studies, we employ two public benchmarks, DiLiGenT Shi et al. (2018) and Luces Logothetis et al. (2022), and our synthetic PS-Verse Testdata. This set contains 441 scenes held out from the training data, with each scene featuring a single object to allow for clearer qualitative evaluation. To verify the model’s generalization capabilities, we test on real-world images from SDM UniPS Ikehata (2023).
4.1 Ablation Study and Comparison Experiment
Table 2: Main ablation with 20 multi-lights input images. The evaluation metrics were measured on our PS-Verse Testdata.
[…]
Uni MS-PS PS-Uni MS-PS 0.72 0.70 9.02
Uni MS-PS PS-Mix 0.63 0.66 10.02
Uni MS-PS PS-Verse 0.75 0.73 7.82
Baseline PS-Verse 0.71 0.69 8.73
+ light register tokens PS-Verse 0.74 0.73 8.13 (0.60 $\downarrow$)
+ global attention PS-Verse 0.80 0.78 6.44 (2.29 $\downarrow$)
+ light alignment PS-Verse 0.86 0.82 5.58 (3.15 $\downarrow$)
+ wavelet branch PS-Verse 0.85 0.82 5.15 (3.58 $\downarrow$)
+ grad perception loss PS-Verse 0.86 0.83 4.84 (3.89 $\downarrow$)
+ curriculum learning PS-Verse 0.88 0.86 4.51 (4.22 $\downarrow$)
Ablation: The main ablation results are presented in Tab. 2. First, to demonstrate the superiority of our PS-Verse dataset, we retrain the Uni MS-PS Hardy et al. (2024) on both the PS-Mix Ikehata (2023) and our PS-Verse. We then compare these models against the officially released Uni MS-PS, which was trained on its native data (PS-Uni MS-PS). The results show that the model trained on PS-Verse yields substantially better feature similarity scores (CSIM and SSIM) and a lower MAE for normal reconstruction, which strongly indicates that PS-Verse is more effective for training high-performance models. Second, we conduct ablations of the various modules within our LINO UniPS method, starting from a baseline model where all our proposed enhancements are removed. Our initial finding is that by merely incorporating unsupervised Light Register Tokens, both the feature similarity metrics and the normal reconstruction results show an improvement. This suggests that these additional tokens can often capture global lighting information even without direct supervision, thereby aiding the disentanglement of normals from illumination. Progressively incorporating our global cross-image attention mechanism to our interleaved Attention Block and light alignment supervision leads to more significant improvements in both feature similarity and normal reconstruction performance. This further validates our central conclusion: that more effectively decoupling lighting and normal features enhances the quality of normal reconstruction. Next, we integrate the wavelet branch to form our Wavelet-based Dual-branch Architecture and the Normal-gradient Perception Loss for extracting fine-grained context. While these two modules have a modest impact on feature similarity metrics, they substantially improve performance on data with complex geometries as intended. Finally, from the last row, adopting curriculum learning boosts all metrics, improving feature similarity while also reducing MAE, which confirms the effectiveness of this training strategy.
Comparison 1: To isolate the architectural advantages of LINO UniPS from the effects of training data, we retrain it alongside Uni MS-PS Hardy et al. (2024) on the PS-Mix dataset Ikehata (2023) to convergence ( 100 epochs). As presented in Tab. 3, our LINO UniPS surpasses SDM UniPS and Uni MS-PS across both evaluated public datasets. Notably, LINO UniPS yields higher CSIM and SSIM scores, quantitatively demonstrating that features extracted by our encoder exhibit greater similarity compared to those from SDM UniPS Ikehata (2023) and Uni MS-PS. It is noteworthy that our LINO UniPS utilizes a decoder architecture identical to that of SDM UniPS Ikehata (2023). This architectural commonality strongly suggests that the observed performance improvements are primarily attributable to our encoder’s enhanced capability to more effectively decouple illumination from geometry. Such effective decoupling fosters stronger normal consistency within the learned features, consequently boosting the accuracy and capability of the final normal reconstruction.
Table 3: Architectural comparison on the PS-Mix dataset Ikehata (2023). The consistency of the encoder features is measured by Cosine Similarity (CSIM) and Structural Similarity (SSIM), while the final normal reconstruction accuracy is evaluated by Mean Angular Error (MAE).
[…]
DiLiGenT Luces
SDM UniPS Ikehata (2023) 0.84 0.72 5.80 13.50
Uni MS-PS Hardy et al. (2024) 0.82 0.76 5.75 13.71
[…]
Comparison 2: We present a comprehensive analysis of the parameter count in Tab. 6. To investigate the impact of model size, we finetune two smaller variants on PS-Verse, Ours-S1 (73.2M) and Ours-S2 (60.4M), by reducing the number of layers in Feature Extractor and Interleaved Attention blocks. These models are designed to have parameter counts comparable to Uni MS-PS (75.5M) and SDM UniPS (59.9M). The results clearly show that our models consistently outperform their counterparts at similar parameter scales, confirming that the superiority of our method stems from its architectural design, not merely from a larger parameter count.
[…]
Table 4: Evaluation on DiLiGenTShi et al. (2018). Uses all 96 images unless otherwise noted (K).
Method Ball Bear Buddha Cat Cow Goblet Harvest Pot1 Pot2 Reading Avg.
UniPS Ikehata (2022) 4.90 9.10 19.40 13.00 11.60 24.20 25.20 10.80 9.90 18.80 14.70
SDM-UniPS Ikehata (2023) 1.50 3.60 7.50 5.40 4.50 8.50 10.20 4.70 4.10 8.20 5.80
Uni MS-PS Hardy et al. (2024) 1.92 3.14 6.16 3.60 4.04 6.35 8.84 4.08 4.88 7.09 5.01
@@ 4.1 Ablation Study and Comparison Experiment
Table 5: Evaluation on LUCES Mecca et al. (2021). Uses all 52 images unless otherwise noted (K).
UniPS Ikehata (2022) 11.01 24.12 23.84 27.90 23.51 28.64 16.24 21.41 35.93 14.53 32.87 28.36 25.36 19.03 23.77
SDM-UniPS Ikehata (2023) 13.30 12.76 8.44 18.58 8.53 19.67 7.25 8.86 26.07 8.30 12.67 15.97 16.01 12.54 13.50
Uni MS-PS Hardy et al. (2024) 10.20 10.52 6.98 12.83 9.60 13.68 6.19 8.33 25.29 6.30 11.47 12.45 11.36 11.79 11.21
@@ 4.2 Quantitative Results
As observed in Tab. 4, on the DiLiGenT Shi et al. (2018), our method achieves a new SOTA performance with an average MAE of 4.65°, outperforming the previous best method. Similar results can also be seen in Tab. 5. On Luces Mecca et al. (2021), a dataset featuring high-frequency information, our method obtains an average MAE of 9.43°. This represents a substantial improvement over the prior best performance (11.21°). Some qualitative results are presented in Fig. 5.
To further underscore the efficacy of our encoder, we conduct an additional experiment. In this setup, we replace the standard decoder in our LINO UniPS with a simpler Multi-Layer Perceptron (MLP) and then finetune this variant until convergence. Our findings are twofold: First, the consistency of the extracted features is further enhanced, as detailed in Fig. 5. Second, although the reconstruction performance sees a slight degradation compared to LINO UniPS with its original, more sophisticated decoder, this MLP-decoder variant (w/mlp) still outperforms SDM UniPS. These results align with our hypothesis that the superior performance of LINO UniPS is primarily driven by its encoder’s advanced capability to generate highly consistent and well-disentangled features.
Moreover, as detailed in Tab. 4 and Tab. 5, our method achieves SOTA on all scenes except for the ’Ball’ object in both benchmarks, demonstrating the general superiority of our approach. Interestingly, on the ’Ball’ scenes, the best performance is achieved by our MLP-decoder variant. We hypothesize that this is because a simpler MLP decoder is better suited for recovering geometrically simple primitives like the ’Ball’.
Table 6: Ablation study on parameter count. Inference time is measured on 16 images (K=16). The $4000\times 4000$ and $512\times 612$ resolution images are sourced from the SDM UniPS real data Ikehata (2023) and the DiLiGenT Shi et al. (2018).
[…]
$4000\times 4000$ $512\times 612$ DiLiGenT LUCES
Ours 84.2 85.1 1.7 4.65 9.43
Ours-S1 73.2 82.9 1.5 4.83 10.05
Ours-S2 60.4 81.0 1.5 4.95 10.89
SDM UniPS 59.9 92.7 1.8 5.83 13.52
Uni MS-PS 75.5 3012.2 35.3 5.01 11.21
Figure 6: Qualitative comparison for the high-resolution (4K) ’Coins and keyboard’ scene from SDM UniPS Ikehata (2023) with 8 input images. Our method recovers more intricate details than the baseline.
@@ 4.3 Qualitative Results
Fig. 1 and Fig. 6 showcase our method’s ability to reconstruct highly detailed and accurate surface normals for real-world objects and scenes, highlighting its strong generalization capabilities.
As shown in Fig. 1, when processing high-resolution images, a close inspection of detail-rich regions like the rabbit’s ears and abdomen reveals a clear distinction. While UniPS and SDM UniPS produce over-smoothed results and fail to capture the intricate surface geometry, our LINO UniPS successfully reconstructs fine-grained details in these challenging areas. Fig. 6 further demonstrates our method’s performance on more complex, high-resolution (4K) real-world scenes. Our method consistently produces more physically plausible and detailed results. For instance, a key differentiator is revealed in the background textile, as shown in the magnified view of the top row. Here, Uni MS-PS, SDM UniPS and UniPS fail to reconstruct the complex texture of the tablecloth, whereas our method demonstrates stronger generalization by accurately recovering its intricate fabric pattern. Additional reconstruction results on real-world scenarios can be found in appendix A1.5.2.
@@ 5 Conclusion
In this paper, we propose LINO UniPS, a novel framework for Universal Photometric Stereo that addresses two core challenges. The first is the failure to disentangle illumination-invariant surface normals from spatially-varying lighting. To this end, we employ Light Register Tokens with an explicit light alignment and an Interleaved Attention Block with a global cross-image attention mechanism. These components work in concert to capture the global lighting context and enable better separation of lighting and normal features. The second challenge is the loss of fine-grained detail. To address this, we integrate a wavelet branch to form a Wavelet-based Dual-branch Architecture and introduce a Normal-gradient Perception Loss, which heightens the model’s sensitivity to intricate geometry. Finally, we contribute a complex and large-scale photometric stereo dataset, hoping to provide valuable reference for future research.
Ethics Statement.
This work does not involve human subjects, personally identifiable information, or sensitive data. The datasets used in this study are publicly available and widely adopted in the machine learning community. All experiments were conducted using standard computational resources without environmental or societal harm. The methodology does not introduce discriminatory biases, and the model’s potential applications are aligned with responsible AI principles. The authors have reviewed the ICLR Code of Ethics and confirm that this submission adheres to its guidelines.
Reproducibility Statement.
To support reproducibility, we provide a complete description of our model architecture, training procedures, hyperparameters, and evaluation protocols in the main paper. Additional implementation details are included in the Appendix. We have strived to document all necessary components with sufficient clarity to enable independent replication of our results.
@@
This supplementary document offers further technical details, demonstrations of the datasets employed, additional insights, and comprehensive results pertaining to our LINO UniPS.
@@
To provide readers with a more in-depth understanding of our LINO UniPS network architecture, in the following, we detail the architecture of our LINO UniPS, breaking it down into its seven key modules: (a) Wavelet DownSample Module, (b) Feature Extractor, (c) Light Registered Attention Module, (d) Fusion Block, (e) Wavelet UpSample Module, (f) Decoder and (g) the training loss. An overview of our network architecture is presented in Fig. 2. In the subsequent sections, we will provide a detailed account of the network’s structural components and the specific transformations of tensor shapes as data progresses through the model.
A1.1.1 Wavelet DownSample Module:
To enable our encoder to extract more fine-grained contexts, we first incorporate the wavelet transform Daubechies (1990), chosen for its ability to separate an image’s high- and low-frequency components Finder et al. (2024); Huang et al. (2017) while concurrently mitigating losses typically incurred during downsampling Peng et al. (2024).
@@
Following SDM UniPS Ikehata (2023), for each pre-processed input image $I_{f}\in\mathbb{R}^{H\times W\times 3}$, we perform two separate transformations: naive downsampling to obtain $I_{f}^{d}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$ in the image domain, and a wavelet transform to yield its corresponding wavelet domain components $I_{f}^{w}\in\mathbb{R}^{4\times\frac{H}{2}\times\frac{W}{2}\times 3}$, namely $I_{f}^{ll}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$, $I_{f}^{lh}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$, $I_{f}^{hl}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$, and $I_{f}^{hh}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$.
A1.1.2 Feature Extractor:
@@
Subsequently, both the downsample image representation $I_{f}^{d}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$ and the set of wavelet components $I_{f}^{w}$ (comprising $I_{f}^{ll},I_{f}^{lh},I_{f}^{hl},I_{f}^{hh}$, each in $\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$) are individually processed. First, each of these input components is partitioned into a sequence of patch-based tokens. These token sequences are then fed into our feature extraction backbone. This ViT backbone is trained during our training procedure to extract rich visual representations from these diverse inputs. In our specific implementation, we set the patch size to $P=8$. Consequently, for each input component with spatial dimensions $H/2\times W/2$, the resulting sequence length is $L=\frac{(H/2)\times(W/2)}{P^{2}}$ tokens. The feature embedding dimension is $D=384$. Following processing by this backbone, we respectively obtain shallow visual feature representations $F_{s,f}^{d}\in\mathbb{R}^{L\times D}$ from the downsample image stream (derived from $I_{f}^{d}$) and $F_{s,f}^{w}\in\mathbb{R}^{4\times L\times D}$ from the wavelet components stream (derived from $I_{f}^{w}$).
A1.1.3 Light Registered Attention Module:
@@
To achieve a more effective decoupling of lighting and normal features that are subsequently processed by the decoder, we introduce our Light Registered Attention Module.